Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Disks

Expert replies
by yourshail123 » Sat Oct 20, 2012 1:24 pm
Each of 4 bags contains 25 blue disks, 25 green disks, 25 orange disks, 25 yellow disks, and nothing else. If one disk is chosen at random from each of the four bags, what is the probability that the number of blue disks chosen will be no less than 1 and no greater than 3?
A) 1/16
B) 41/128
C) 87/128
D) 225/256
E) 255/256
Join the discussion
Source: — Problem Solving |

by TheGmatTutor » Sat Oct 20, 2012 2:51 pm
I'd rephrase the question as "What is the probability you will pick 1, 2, or 3 blue disks?" In order to answer this question, I would start with 1 and subtract off the probability of zero blue disks and the probability of 4 blue disks.

1 - P(0 blue disks) - P(4 blue disks)

1 - (3/4)^4 - (1/4)^4

1 - (81/256) - (1/256)

1 - (82/256)

174/256

87/128

Answer C
If this post was helpful, please click the "Thanks" button.
Join the discussion

by yourshail123 » Sun Oct 21, 2012 9:39 am
I am expecting some one to try with the conventional way - Calculating out the particular probability for picking 1, 2, or 3 blue disks. Anyone?
Agree with TheGmatTutor - your way is best way possible.
Join the discussion

by beingNeha » Sat Jul 06, 2013 10:23 pm
Probability of getting 0 blue disks p(0) = 3/4 * 3/4 * 3/4 * 3/4 = 81/256
Probability of getting 4 blue disks p(4) = 1/4 * 1/4 * 1/4 * 1/4 = 1/256

Probability of getting 1,2, or 3 blue disks p(1, 2, 3) = 1 - 81/256 - 1/256 = 174/256 = 87/128 (C). B-)
Join the discussion

by Uva@90 » Mon Jul 08, 2013 8:19 pm
Hi All,
Could any one please explain me how probability of zero blue disk is (3/4)^4 and probability of 4 blue disk is (1/4)^4.

Thanks in advance.
Regards,
Uva.
Join the discussion

by Jeff@TargetTestPrep » Wed Nov 08, 2017 4:40 pm
yourshail123 wrote:Each of 4 bags contains 25 blue disks, 25 green disks, 25 orange disks, 25 yellow disks, and nothing else. If one disk is chosen at random from each of the four bags, what is the probability that the number of blue disks chosen will be no less than 1 and no greater than 3?
A) 1/16
B) 41/128
C) 87/128
D) 225/256
E) 255/256

We should first note that the condition “the number of blue disks chosen is no less than 1 and no greater than 3” is equivalent to the condition “the number of blue disks is neither 0 nor 4.”

We can use the following equation:

P(the number of blue disks chosen will be no less than 1 and no greater than 3) = 1 - P(selecting 0 blue disks) - P(selecting 4 blue disks)

Let’s first determine P(selecting 0 blue disks):

75/100 x 75/100 x 75/100 x 75/100 = (3/4)^4 = 81/256

Next let’s determine P(selecting 4 blue disks):

25/100 x 25/100 x 25/100 x 25/100 = (1/4)^4 = 1/256

Thus:

P(the number of blue disks chosen will be no less than 1 and no greater than 3) is:

1 - 81/256 - 1/256 = 1 - 82/256 = 174/256 = 87/128

Answer: C

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion

by Matt@VeritasPrep » Thu Nov 09, 2017 7:06 pm
Uva@90 wrote:Hi All,
Could any one please explain me how probability of zero blue disk is (3/4)^4 and probability of 4 blue disk is (1/4)^4.

Thanks in advance.
Regards,
Uva.
p(zero blue disks) = p(first disk isn't blue) * p(second disk isn't blue) * p(third disk isn't blue) * p(fourth disk isn't blue)

In any given bag, the ratio of blues to nonblues is 25 : 75, so the probability of getting a nonblue disk is 75/(75 + 25), or 3/4. With that in mind, p(zero blue) = 3/4 * 3/4 * 3/4 * 3/4.
Join the discussion